OpenStudy (anonymous):

solving quadratic equations by graphing and very confused!!! please help!! y=x^2+5x=4

OpenStudy (anonymous):

i mean y=x^+5x+4 not equals sorry :)

hero (hero):

y = x^2+5x+4 y = x^2+4x + x + 4 y = x(x+4)+1(x+4) y = (x+4)(x+1)

OpenStudy (anonymous):

x = -4 or -1

OpenStudy (anonymous):

still need help

hero (hero):

With?

OpenStudy (anonymous):

solving for vertex im not getting the right answer

hero (hero):

x = -b/(2a)

hero (hero):

x= -5/2 = -2.5

hero (hero):

y = 6.25 -12.5 +4

OpenStudy (anonymous):